What volume of 0.817 M KOH solution is required to make 3.68 L of a solution with pH of 12.6?
Answer #1

let required volume be V in L
number of mole of KOH = (molarity)*V
= 0.817V
molarity of new solution = (number of mole)/(new volume in L)
=0.817V/3.68
= 0.222V

since, KOH is strong base
[OH-] = [KOH]
[OH-] = 0.222V

pOH = -log[OH-]
14 - pH = -log(0.222V)
14-12.6 = -log(0.222V)
-1.4 = log(0.222V)
0.222V = 10^-1.4
0.222V = 0.0398
V = 0.179 L
= 179 mL

Answer : 179 mL